Fermi National Laboratory


Accelerator Update for August 3, 2005 - August 5, 2005

Wednesday August 3
The day shift began with Operations monitoring store 4315, with the Recycler stashing antiprotons, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using colliding beam, with MiniBooNE and NuMI taking beam, and with no beam to the Meson area due to stashing.

The Recycler ended its stashing and the Antiproton Source (Pbar) resumed stacking at 8:55 AM.

A Main Injector (MI) expert began a slipstacking study at 9:53 AM.

At 10:56 AM, Operations reported that a MI RF station (MIRF1) tripped off. RF experts said they needed to change the modulator. They finished this work and stacking resumed at 11:33 AM.

MI experts began a three-hour MI-8 transfer line beam-matching study at 12:08 PM.

At 12:35 PM, Operations reported a Booster kicker (MKS04) tripped off due to sparking. Experts investigated.

At 1:30 PM, MIRF1 tripped off and wouldn't reset. Experts investigated and declared its power amplifier had failed. They bypassed the station.

Booster experts reported at 1:41 PM that they had reterminated a cable for a MKS04.

Operations put the Recycler into shot setup at 4:22 PM to stash antiprotons.

At 4:46 PM, Operations reported that NuMI had tripped off for the first of eight times in about five hours. It reset each time.

The Recycler stash ended and Pbar resumed stacking at 5:38 PM.

Thursday August 4
The midnight shift began with Operations monitoring store 4315 and the Recycler stash, with Pbar stacking, and with MiniBooNE, NuMI, MTest, and MCenter taking beam.

MTest experimenters said at 12:07 AM that they were done with the beam for the night.

Tevatron (TeV) experts began a study at 3:03 AM.

Operations terminated store 4315 at 5:13 AM and began preparing TeV sector F for access.

The TeV access ended and Operations began establishing beam to MCenter and MTest at 9:46 AM.

Pbar experts began making transfer line lattice measurements at 10:52 AM. They finished their study at 11:27 AM.

Operations put the TeV through a dry squeeze at 12:31 PM.

Operations put the TeV through a wet squeeze at 1:47 PM.

Operations put the TeV into shot setup at 2:09 PM, but ten minutes later a BPM problem put the shot setup on hold.

Experts resolved the BPM problem and shot setup continued at 3:19 PM.

Operations established combination store 4318 at 5:24 PM with an initial luminosity of 129.8E30. This set a NEW LUMINOSITY RECORD. Pbar resumed stacking.

At 5:28 PM, Operations reported that NuMI was experiencing beam problems. After a lot of investigating and contacting experts, Operations determined that the problem was with slipstacking. MI experts were called in.

Operations put the Recycler into shot setup at 9:59 PM. The stash ended and Pbar resumed stacking at 10:43 PM.

At 11 PM, a NuMI power supply (E:HV101) tripped off and wouldn't reset. An operator called in an expert.

Friday August 5
The midnight shift began with Operations monitoring store 4318 and stash, with Pbar stacking, with NuMI off due to power supply problems, and with MiniBooNE and MCenter taking beam.

The power supply expert reported at 1:01 AM that he had replaced an AC relay breaker. NuMI resumed taking beam.

MI experts reported at 1:18 AM that they had resolved the slipstacking problem.

The Recycler began stashing antiprotons at 3:40 AM. The stash ended and Pbar and MCenter resumed taking beam.

At 6:26 AM, the output for a Booster RF station (BRF13) dropped to zero and that caused several other stations to start sparking. An operator called in an expert. At 7:29 AM, he reported the problem was in the tunnel. BRF13 was bypassed.

At 7:53 AM, a problem developed and cleared for a Linac quadrupole power supply (QPS113). Later, an expert changed out a control card for this power supply.

The Plans for Friday and the Weekend
The plan for today is to allow some machine studies. The plans for the weekend are to stack, stash, and store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 8/5/2005   email Fermilab